How fast must a softball travel to reach 35 feet in 1.20 seconds?
This formula will greatly help me to know how fast my pitchers must pitch the ball so my catchers can make the throws to catch runners stealing second base.

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

How fast must a softball travel to reach 35 feet in 1.20 seconds?
This formula will greatly help me to know how fast my pitchers must pitch the ball so my catchers can make the throws to catch runners stealing second base.
-------------
If you mean the speed to throw it to do that:
d = r*t
r = d/t = 35/1.2 = 29.2 ft/sec
= 19.89 mi/hr
---------------
It has to be faster than that, tho, because the ball won't travel in a straight line.

Answer by stanbon(75887)

How fast must a softball travel to reach 35 feet in 1.20 seconds?
------
[(35/1.2) (ft/sec)]
---
= (35/5280)/(1.2/3600)](miles/hour)
---
= 19.89 mph
===============
Cheers,
Stan H.